OpenStudy (driftracer305):

Inverse of (2x-3)^2............pls help

OpenStudy (akashdeepdeb):

Let\[y = (2x-3)^2\]\[\sqrt{y} = \pm (2x-3)\]\[\pm \sqrt{y} + 3 = 2x\]\[x = \frac{\pm \sqrt{y} + 3}{2}\] Now since an invertible function must be one-one and onto that is every y image must have a Unique x pre-image, we take only either + or -. So finally, \[x = \frac{\sqrt{y} + 3}{2}\] Getting this?

OpenStudy (anonymous):

To obtain the inverse function, you just have to replace y by x and x by y\[y=(2x-3)^2\]replacing\[x=(2y-3)^2\]now we have to isolate the y again\[\sqrt{x}=2y-3\]then we got\[y=\frac{\sqrt{x}+3}{2}\]

OpenStudy (akashdeepdeb):

Another thing, what I wrote there is \(g(y)\) and not \(f^{-1} (x)\). @D3xt3R has explained it there! ^

OpenStudy (driftracer305):

ok, thanks a lot guys @AkashdeepDeb and @D3xt3R .....

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

;D